OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? r + (-12) = 19

OpenStudy (anonymous):

12 should be added to both sides t solve for the variable r + (-12 + 12) = 19 + (12) r + 0 = 31 r = 31

OpenStudy (texaschic101):

I agree....add 12 to both sides